A rod 2m long and of diameter 30mm is subjected to an axial pull of 30kN. If the young’s
modulus of the material of the rod is 2x105 N/mm2
, determine
a) stress
b) strain and
c) Elongation
1
Expert's answer
2020-09-07T08:17:48-0400

the cross sectional area,A = (22/7d2)/4= (22/7*302)/4 = 707.143mm2

a)stress = load/area = 30000/707.143 = 42.424N/mm2

b)strain = stress/young's modulus = 42.424/200000= 0.00021212

c)elongation = load*length/area* young's modulus = 30000*2000/707.143*200000 = 0.424mm